# Jammer Mitigation in Absorptive RIS-Assisted Uplink NOMA

### v1

@@ Abstract
Non-orthogonal multiple access (NOMA) is a promising technology for next-generation wireless communication systems due to its enhanced spectral efficiency. In this paper, we consider an uplink NOMA system operating together with a high-dimensional absorptive reconfigurable intelligent surface (A-RIS). We aim to minimize the total power transmitted by the users in order to meet signal-to-interference-plus-noise constraints at the base station in the presence of a jammer. We propose an iterative algorithm to solve the high-dimensional non-convex optimization problem using linear programming to find the transmit powers and a fractional programming algorithm based on the Dinkelbach algorithm with a sequential convex relaxation procedure to optimize the reflection coefficients. We show that our algorithm converges on large optimization problems, with a jammer comprising as many as $64$ antennas, and an A-RIS with $128$ elements. Our numerical results show that, compared with a standard RIS that reflects all impinging energy, the A-RIS can dramatically reduce the users’ required transmit power and successfully mitigate interference from the jammer. The absorption capability of the A-RIS is in particular useful in cases when the number of jammer antennas is of the same order as the number of A-RIS elements.
@@
A-RIS, NOMA, phase shift, non-convex optimization, SDR, Dinkelbach algorithm, absorption, interference, jammer.
@@ I Introduction
Non-orthogonal multiple access (NOMA) is a promising technique for next-generation radio access. In contrast to conventional orthogonal multiple access (OMA) schemes, where each user is served using a single orthogonal resource block, NOMA is of particular interest because it allows multiple users to share the same orthogonal time, frequency, spatial, and code-domain resource blocks. For OMA, there is no interference between multiple users since they are assigned to orthogonal resource blocks, but conventional OMA schemes are restricted in terms of the maximum number of supported users, due to the limited availability of orthogonal resources. NOMA has significant advantages over traditional OMA schemes, with higher effective bandwidth, support for massive connectivity, and reduced transmission latency. By enabling multiple users to share resources, NOMA can more efficiently handle a higher number of users[1].
Generally, NOMA is divided into two main classes: power-domain and code-domain [2]. Power-domain NOMA exploits situations where the users have different path loss levels. In downlink NOMA this implies that the users nearer to the base station (BS) typically have better channel conditions compared to distant users, who require higher transmission power to mitigate the higher path loss. The idea behind power-domain NOMA is that the users nearer the BS can employ successive interference cancellation (SIC) to remove the strong signals destined for the remote users before decoding their own signal[3].
@@
In contrast to active amplify-and-forward relays that usually operate in half-duplex mode, the RIS functions in full-duplex mode and only reflects the received signals as a passive array, without the expenditure of any radio-frequency energy, which enables spectrum-efficient and cost-effective communications [7], [8], [9]. An RIS can increase the strength of received signals through beamforming gain and constructive interference, eliminate interfering signals through destructive interference or null-steering, and provide virtual line-of-sight paths to overcome blockages. For single-antenna NOMA systems in the presence of a jammer, an RIS can provide a large number of spatial DoFs that allow the network to cancel the jamming signal, essentially creating a null channel from the jammer to either the single-antenna BS or single-antenna users. This is a key capability that will be studied in this paper.
In a conventional RIS implementation, it is the phase shift of the reflection coefficient of each element that is adjusted in order to achieve the desired effect on the wireless channels. More recently, researchers have also studied RIS architectures where both the phase shift and the attenuation or absorption of the reflection coefficient of each element can be individually controlled, referred to as absorptive RIS (A-RIS) [10]. In this latter case, the energy absorbed by the A-RIS can be refracted to directions on the other side of the surface [11], [12] sampled using an active RF receiver for channel estimation or sensing [13], [14],[15] or simply dissipated. In [16], the interference mitigation problem using A-RIS in two spectrum sharing scenarios is investigated: under spectral coexistence of radar and communication systems, and spectrum sharing in device-to-device (D2D) communications. The results show that A-RIS significantly outperforms non-absorptive RIS for interference suppression scenarios. Similarly, the interference cancellation ability of A-RIS was exploited to support joint D2D and cellular communications in[17].
RIS technology has been applied in many different types of wireless communications scenarios, including NOMA. However, while the use of conventional phase-shift-only RIS has been proposed for NOMA applications with the goal of improving spectral efficiency, to the best of our knowledge there is no prior work reported on using A-RIS with NOMA, nor on using RIS to mitigate the impact of external interference (e.g. jamming) on NOMA performance. NOMA is a vulnerable technology due to the use of shared resources since the decoding performance depends on sufficient differences in the users’ received power levels, which in the uplink involves adapting the transmit power levels to the channel conditions. However, jamming can reduce the efficiency of NOMA due to the severe loss of signal-to-noise ratio (SNR) [18].
@@ I-A Related Works
Several recent works have studied the use of RIS in NOMA applications. For example, an RIS-assisted NOMA system was investigated in [19] to achieve a tradeoff between sum rate and power consumption. The results indicate that RIS-assisted NOMA can achieve better energy efficiency performance than the OMA scheme. In [20], joint power allocation, hybrid beamforming, and phase shift optimization for downlink multi-user RIS-assisted mm-wave-NOMA was designed to maximize the sum rate using alternating manifold optimization (AMO) and successive convex approximation (SCA) methods. The analytical results reveal that the achievable sum rate was significantly improved by employing optimized RIS phase shifts. In [21], the maximum sum rate achieved by optimizing the RIS phase shifts subject to the user power constraints was studied for NOMA. The results show that the proposed approaches improve spectral efficiency through the use of different power levels. In [22], NOMA was implemented in a multi-cell scenario assisted by multiple RIS to minimize the transmit power in the uplink. The results demonstrate that inter-group interference cancellation in NOMA, with the help of multi-reflection RIS, achieves lower total transmit power. An RIS-aided multi-unmanned aerial vehicle (UAV) system with NOMA was studied in [23], with the goal of minimizing the system’s power consumption. The simulation results demonstrate that combining RIS with multi-UAV-assisted NOMA networks significantly reduces the total power consumption. In [24], RIS-aided multi-cell NOMA was shown to improve the coverage probability using stochastic geometry methods and two different channel models. The analytical results show that increasing the number of RIS elements can improve performance and provide extra flexibility for NOMA decoding orders. Furthermore, many other research problems in the context of RIS-assisted NOMA have been studied using different optimization methods in [25], [26].
Relatively little work has been done on optimizing NOMA performance in the presence of a jammer. In [27] optimal user grouping for NOMA was proposed to overcome the impact of jamming and improve the sum rate. In [28], a mobile access point or a UAV was exploited together with joint power control to mitigate the effect of a jamming attack and increase the reliability of the communication. Furthermore, anti-jamming precoding was proposed in [29] to minimize the total transmit power in an uplink MIMO-NOMA system. Finally, in [30], transmit beamforming together with the use of an RIS and artificial noise was proposed to enhance the secrecy of a NOMA system. However, except our initial work in [25], none of the previous works cited above have considered the use of an RIS for mitigating jamming for a NOMA system, either in the uplink or downlink.
@@ I-B Motivations and Contributions
As already mentioned above, the focus of this paper is on the use of RIS, and in particular A-RIS, to mitigate the influence of a jammer on an uplink NOMA system. Our objective is to minimize the sum transmit power of the uplink NOMA users under quality of service constraints in the presence of a jammer. This problem formulation is well aligned with 6G key requirements on sustainability, trustworthiness, and digital inclusion [31]. The main contributions of our work are summarized as follows:
  • We consider an uplink power domain NOMA system assisted by an A-RIS in order to mitigate interference from a multi-antenna jammer.
  • We propose an optimization framework that minimizes the total transmit power of the user terminals under constraints on the A-RIS response and the individual signal-to-interference-plus-noise ratio (SINR) at the BS for each of the users. This turns out to be a highly complex non-convex optimization problem. To address it, we propose a solution based on iterating between Linear Programming (LP) to find the user powers and sequential convex relaxation based on Dinkelbach’s algorithm [32] to find the A-RIS response.
@@
Organization: The rest of the paper includes the following sections. In Section II we detail the targeted scenario and the A-RIS-assisted NOMA system model, and we define the resulting optimization problem. In Section III, we provide details on the implementation of the A-RIS which motivates our system model and simulation assumptions. Section IV then provides details for our proposed optimization algorithm that alternates between LP and the Dinkelbach algorithm combined with semi-definite relaxation (SDR) to minimize the total power transmitted from the users to the BS. Section V then presents numerical results that validate the performance of the algorithm in challenging scenarios. In Section VI we then elaborate on the relevance of the results for the downlink scenario and on additional practical A-RIS considerations for future work. Finally, we summarize our findings in Section VII.
@@ II System Model and Problem Formulation
In this section, we first describe the system and signal model for our considered A-RIS-aided NOMA system, and provide assumptions regarding the channel state information (CSI) available to the BS and the jammer.
@@ II-A System Model
Assuming the $K$ users are indexed in the order of declining channel gains, and $p_{1}|h_{1}|^{2}\geq p_{2}|h_{2}|^{2}\geq...\geq p_{k}|h_{k}|^{2}$, consistent with the situation illustrated in Fig. 1. The BS then first decodes the signal for UE1 treating the multi-user interference of the signal from the other users as noise, and then subtracts it from the received signal $r$ when decoding the signal from UE2. Thus, the SINR for decoding UE1 is given by
@@
Continuing on in this fashion and assuming perfect SIC for UE2 and each of the subsequent user’s signals with no error propagation, the SINR for the last user, UEK can be written as
@@ II-B Jammer Modeling
[…]
@@ III Absorptive RIS
In Section V we show that the ability to adjust the magnitude of the reflected power at each RIS element provides additional DoFs that are particularly useful in situations like the one we consider in this work, in which interference mitigation is required.
@@ III-B Practical Design of A-RIS
Compared to widely studied phase-controlled RIS designs, A-RIS designs with concurrent amplitude and phase reconfigurability are less prevalent due to the challenges in physical implementation and availability of tunable components [33]. Nevertheless, in recent years, the higher flexibility and versatility provided by combined amplitude-phase control has sparked growing research interest in this technology in the antenna, optics, and materials communities [34]. As an example, take the reconfiguration mechanism based on using semiconductor p-i-n and varactor diodes. A p-i-n diode can not only function as a microwave switch to generate a discrete phase shift but can also operate as a variable resistor by applying a forward current to bias the diode’s junction to the transition state. In this case, the junction resistance can vary by several orders of magnitude [34]. When an RIS unit cell (UC) combines a p-i-n diode with a varactor diode, exploiting the control of the reverse-biased junction capacitance, a variable amplitude-phase RIS reflection response can be engineered. Recently, such a reconfigurable RIS UC was experimentally demonstrated, allowing independent and continuous amplitude-phase control.
@@ IV-A Problem Formulation for $K$ NOMA Uplink Users
There are several ways to illustrate the benefit of the A-RIS in the problem under consideration. Here we study the problem of minimizing the total transmit power of the users such that given SINR quality-of-service constraints of the users are met in the presence of a jammer. The optimization problem for $K$ users can be written as follows:
@@ IV-B Proposed Solution
[…]
@@
[…]
@@ IV-C Algorithm Initialization
[…]
@@ V-A Performance for Various Number of Users with Fixed A-RIS Location
Here we study the performance of the proposed algorithm in terms of the sum total uplink transmit power for $K=2$ and $K=3$ users, for various numbers of jammer antennas and A-RIS elements, assuming a fixed location of the A-RIS. We also assess the benefit and characterize the optimized behavior of the absorption capability of the A-RIS in terms of the average reflection coefficients of the A-RIS elements. In Table I we show our simulation parameters.
@@
We choose to focus on a jammer with a rather large number of antennas $M$ and also an A-RIS with a large number of elements $N$. This is an interesting case in practice, since in mm-wave systems a rather large number of antennas/elements would be needed, and we want to show the capability of Algorithm 1 to find a solution for such large optimization problems. This is in contrast to our initial work in [25] using the Matlab fmincon tool, in which we could not find an optimal solution for a jammer larger than $M=16$ in combination with A-RIS larger than $N=32$. Note that we assume a maximum A-RIS reflection coefficient $\beta_{\text{max}}=1$, despite our discussions in Section III. The reason is that in our evaluation scenarios a value of $\beta_{\text{max}}<1$ can simply be included in the assumed path losses to the A-RIS.
@@
In Fig. 4 we show the required total user transmit power to meet the user quality of service requirements in the presence of a jammer with $M=4,8,16$ antennas, as a function of the number of A-RIS elements $N=32,38,44,64,88,100,128$ for $K=2$ users. As seen, a large A-RIS is very beneficial to mitigate the jammer’s power. The sum transmit power drops from 0.28 W (24.5 dBm) to 0.07 W (18.5 dBm) when increasing the number of A-RIS elements from $=4$ to $N=128$ in the case of a jammer with $M=4$ antenna elements. Doubling the number of jammer antennas to $M=8$ does not significantly change the required total transmit power. However, with $M=16$ there is a significant increase in the required sum user power for A-RIS with $32$ elements, increasing to 0.54 W (27.3 dBm), whereas for large A-RIS the increase in sum user power is still small.
@@
In Fig. 6 we further investigate the absorption behavior of the A-RIS as a function of the number of jammer antennas $M$ and A-RIS elements $N$. Here we investigate larger $M$ and $N$ compared to Fig. 5, but the behavior of the A-RIS is the same. This also shows that Algorithm 1 converges well also for large $M$ and $N$. In general, for the investigated deployment and propagation scenario here, it seems that with $N/M\geq\approx 5$, absorption is not needed, whereas for $N/M<\approx 5$ absorption is very useful. This can be explained by the required number of DoFs for the A-RIS to cancel the jammer signal.
In Fig. 7 and Fig. 8, we investigate the performance and behavior of the A-RIS in the case of $K=3$ users. Algorithm 1 also converges very well in this case, and the overall behavior is the same as in the $K=2$ users case. One difference comparing Fig. 7 with Fig. 4 is that the total required transmit power by the users is larger for $K=3$ compared to $K=2$, due to the residual NOMA multi-user interference for the users $k<K$ that is not canceled at the BS. Comparing Fig. 6 and Fig. 8 shows that the A-RIS absorption decreases for the larger number of users with increased multiuser interference (MUI). This is interesting since absorption increases when the jammer interference increases. The difference can be explained by the fact that the A-RIS is attempting to entirely cancel the jammer while combating the MUI still requires the individual user signals to be received with strong SNR.
Figure 7: Required total transmit power to meet the user quality of service requirements in the presence of a jammer with $M=4,8,16$ antennas, as a function of the number of A-RIS elements in case of $K=3$ users.

### v2

@@ Abstract
Non-orthogonal multiple access (NOMA) is a promising technology for next-generation wireless communication systems due to its enhanced spectral efficiency. However, wireless communication is facing increasing requirements for security. To that end, jamming mitigation using multi-antennas has emerged as an important research topic. In this paper, we consider an uplink NOMA system with a reconfigurable intelligent surface (RIS) that assists the uplink users and, at the same time, mitigates the jammer. Our goal is to minimize the total users’ transmitted power under signal-to-interference-plus-noise ratio constraints at the base station. To be effective, typically a high-dimensional RIS is needed, leading to a large optimization problem, which in general faces convergence problems. We propose an iterative algorithm for this high-dimensional non-convex optimization problem that converges with a jammer comprising as many as 64 antennas, and an RIS with 128 elements. More specifically, we introduce a design and optimize the performance of an absorptive RIS (A-RIS). Compared to a standard RIS, we show that an A-RIS can dramatically reduce the users’ required transmit power and successfully mitigate the jammer. The A-RIS is in particular useful in cases when the number of jammer antennas is of the same order as the number of A-RIS elements.
@@
A-RIS, NOMA, phase shift, non-convex optimization, SDR, Dinkelbach algorithm, absorptive, interference, jammer.
@@ I Introduction
Non-orthogonal multiple access (NOMA) is a promising technique for next-generation radio access. In contrast to conventional orthogonal multiple access (OMA) schemes, where each user is served using a single orthogonal resource block, NOMA is of particular interest because it allows multiple users to share the same orthogonal time, frequency, spatial, and code-domain resource blocks. NOMA has significant advantages over traditional OMA schemes, with higher effective bandwidth, support for massive connectivity, and reduced transmission latency [1].
Generally, NOMA is divided into two main classes: power-domain and code-domain [2]. Power-domain NOMA exploits situations where the users have different path loss levels. The idea behind power-domain NOMA is that the users nearer the base station (BS) can employ successive interference cancellation (SIC) to remove the strong signals destined for the remote users before decoding their own signal [3].
@@
In contrast to active amplify-and-forward relays that usually operate in half-duplex mode, the RIS functions in full-duplex mode and only reflects the received signals as a passive array, without the expenditure of any radio-frequency energy, which enables spectrum-efficient and cost-effective communications [7], [8], [9]. An RIS can increase the strength of received signals through beamforming gain and constructive interference, eliminate interfering signals through destructive interference or null-steering, and provide virtual line-of-sight paths to overcome blockages. For single-antenna NOMA systems in the presence of a jammer, an RIS has the potential to provide a large number of spatial DoFs that could allow the network to cancel the jamming signal, essentially creating a null channel from the jammer to either the single-antenna BS or single-antenna users. This is the key RIS capability that will be studied in this paper in the context of jamming mitigation in uplink NOMA.
In a conventional RIS implementation, it is the phase shift of the reflection coefficient of each element that is adjusted in order to achieve the desired effect on the wireless channels. More recently, researchers have also studied RIS architectures where both the phase shift and the attenuation or absorption of the reflection coefficient of each element can be controlled individually, referred to as absorptive RIS (A-RIS) [10]. In this latter case, the energy absorbed by the A-RIS can be refracted to directions on the other side of the surface [11], [12] sampled using an active RF receiver for channel estimation or sensing [13], [14],[15] or simply dissipated. In [16], the interference mitigation problem using A-RIS in two spectrum sharing scenarios is investigated: under spectral coexistence of radar and communication systems, and spectrum sharing in device-to-device (D2D) communications. The results show that A-RIS significantly outperforms non-absorptive RIS for interference suppression scenarios. Similarly, the interference cancellation ability of A-RIS was exploited to support joint D2D and cellular communications in [17].
RIS technology has been applied in many different types of wireless communications scenarios, including NOMA. However, while the use of conventional phase-shift-only RIS has been proposed for NOMA applications with the goal of improving spectral efficiency, to the best of our knowledge, there is no prior work reported on using A-RIS with NOMA, nor on using RIS to mitigate the impact of external interference (e.g., jamming) on NOMA performance. NOMA is a vulnerable technology due to the use of shared resources since the decoding performance depends on sufficient differences in the users’ received power levels, which in the uplink involves adapting the transmit power levels to the channel conditions. However, jamming can reduce the efficiency of NOMA due to the severe loss of signal-to-noise ratio (SNR) [18].
@@ I-A Related Works
Several recent works have studied the use of RIS in NOMA applications. For example, an RIS-assisted NOMA system was investigated in [19] to achieve a tradeoff between sum rate and power consumption. In [20], joint power allocation, hybrid beamforming, and phase shift optimization for downlink multi-user RIS-assisted mm-wave NOMA was designed to maximize the sum rate using alternating manifold optimization (AMO) and successive convex approximation (SCA) methods. In [21], the maximum sum rate achieved by optimizing the RIS phase shifts subject to the user power constraints was studied for NOMA. In [22], NOMA was implemented in a multi-cell scenario assisted by multiple RISs to minimize the transmit power in the uplink. An RIS-aided multi-unmanned aerial vehicle (UAV) system with NOMA was studied in [23], with the goal of minimizing the system’s power consumption. In [24], RIS-aided multi-cell NOMA was shown to improve the coverage probability using stochastic geometry methods and two different channel models. Furthermore, many other research problems in the context of RIS-assisted NOMA have been studied using different optimization methods in [25], [26].
Relatively little work has been done on optimizing NOMA performance in the presence of a jammer. In [27], optimal user grouping for NOMA was proposed to overcome the impact of jamming and improve the sum rate. In [28], a mobile access point or a UAV was exploited together with joint power control to mitigate the effect of a jamming attack and increase the reliability of the communication. Furthermore, anti-jamming precoding was proposed in [29] to minimize the total transmit power in an uplink MIMO-NOMA system. In [30], transmit beamforming together with the use of an RIS and artificial noise was proposed to enhance the secrecy of a NOMA system. Several works have studied the use of RIS for anti-jamming in [31],[32], and [33]. However, except for our initial work in [25], none of the previous works cited above have considered the use of an RIS for mitigating jamming for an uplink NOMA system.
@@ I-B Motivations and Contributions
As already mentioned above, the focus of this paper is on the use of RIS, and in particular A-RIS, to mitigate the influence of a jammer on an uplink NOMA system. Our objective is to minimize the sum transmit power of the uplink NOMA users under quality of service (QoS) constraints in the presence of a jammer. This problem formulation is well aligned with 6G key requirements on sustainability, trustworthiness, and digital inclusion [34]. The main contributions of our work are summarized as follows:
  • We show that an uplink power-domain NOMA system assisted by an A-RIS can play a key role in mitigating interference from a sophisticated multi-antenna jammer. Compared to an uplink NOMA system without A-RIS assistance, we show a gain of more than 28 dB in the required sum transmit power by the uplink users. These gains come from the combined capability of an A-RIS to not only mitigate the jammer interference but also enable substantial beamforming and interference coordination gains for the uplink NOMA users. Unlike conventional RIS, we show that the nature of A-RIS introduces additional new important degrees of freedom in optimizing the system performance while keeping power consumption low.
  • We propose a novel joint optimization problem formulation that minimizes the total transmit power of user terminals under constraints on the A-RIS response and the individual signal-to-interference-plus-noise ratio (SINR) at the BS for each of the users. This problem has not been studied in the literature before, and it turns out to be a highly complex non-convex optimization problem due to the strong coupling between user power control and the A-RIS phase-shift and gain design.
  • To efficiently solve this problem, we develop an iterative algorithm that combines Linear Programming (LP) to find the user powers, ensuring optimal power allocation in each iteration and sequential convex relaxation via Dinkelbach’s algorithm [35] to efficiently optimize the A-RIS response, leveraging fractional programming principles. The proposed algorithm does not rely on conventional convex approximations such as SDR (Semidefinite Relaxation), making it more computationally scalable and adaptable to large-scale A-RIS, which is a key enabler for realistic A-RIS deployments. Our approach effectively balances power minimization and interference mitigation.
@@
Organization: The rest of the paper includes the following sections. In Section II, we detail the targeted scenario and the A-RIS-assisted NOMA system model, and we define the resulting optimization problem. In Section III, we provide details on the implementation of the A-RIS, which motivates our system model and simulation assumptions. Section IV then provides details for our proposed optimization algorithm that alternates between LP and the Dinkelbach algorithm combined with semi-definite relaxation (SDR) to minimize the total power transmitted from the users to the BS. Section V then presents numerical results that validate the performance of the algorithm in challenging scenarios, followed by a discussion on the relevance of the results for a corresponding downlink scenario. Finally, we summarize our findings in Section VI.
@@ II System Model and Problem Formulation
In this section, we first describe the system and signal model for our considered A-RIS-aided NOMA system and provide assumptions regarding the channel state information (CSI) available to the BS and the jammer.
@@ II-A System Model
Assume the $K$ users are indexed in the order of declining channel gains, and $p_{1}|h_{1}|^{2}\geq p_{2}|h_{2}|^{2}\geq...\geq p_{k}|h_{k}|^{2}$, consistent with the situation illustrated in Fig. 1. The BS then first decodes the signal for UE1 treating the multi-user interference of the signal from the other users as noise, and then subtracts it from the received signal $r$ when decoding the signal from UE2. Thus, the SINR for decoding UE1 is given by
@@
Continuing to assume a perfect SIC for UE2 and each of the subsequent user’s signals with no error propagation, the SINR for the last user, UEK can be written as
@@ II-B Jammer Modeling
$\mbox{\rm s.t.}\quad E\left(\|{\bf x}_{\text{j}}\|^{2}\right)\leq P_{\text{j}}\;,$
@@ III Absorptive RIS
In Section V, we show that the ability to adjust the magnitude of the reflected power at each RIS element provides additional DoFs that are particularly useful in situations like the one we consider in this work, in which interference mitigation is required.
@@ III-B Practical Design of A-RIS
Compared to widely studied phase-controlled RIS designs, A-RIS designs with concurrent amplitude and phase reconfigurability are less prevalent due to the challenges in physical implementation and availability of tunable components [36]. Nevertheless, in recent years, the higher flexibility and versatility provided by combined amplitude-phase control has sparked growing research interest in this technology in the antenna, optics, and materials communities [38]. As an example, take the reconfiguration mechanism based on using semiconductor p-i-n and varactor diodes. A p-i-n diode can not only function as a microwave switch to generate a discrete phase shift but can also operate as a variable resistor by applying a forward current to bias the diode’s junction to the transition state. In this case, the junction resistance can vary by several orders of magnitude [38]. When an RIS unit cell (UC) combines a p-i-n diode with a varactor diode, exploiting the control of the reverse-biased junction capacitance, a variable amplitude-phase RIS reflection response can be engineered. Recently, such a reconfigurable RIS UC was experimentally demonstrated, allowing independent and continuous amplitude-phase control [37].
@@ IV-A Problem Formulation for $K$ NOMA Uplink Users
There are several ways to illustrate the benefit of the A-RIS in the problem under consideration. Here we study the problem of minimizing the total transmit power of the users such that given SINR QoS constraints of the users are met in the presence of a jammer. The optimization problem for $K$ users can be written as follows:
@@
Note that in our optimization problem formulation above, we ensure robust SIC by incorporating SINR constraints into the optimization problem, which enforce a minimum threshold to guarantee the required QoS for each user. By setting these constraints appropriately, the system automatically achieves the conditions necessary for successful SIC. Therefore, even without explicitly defining SIC constraints, our approach inherently ensures their fulfillment.
@@ IV-B Proposed Solution
$\text{s.t.}~p_{k}\mathop{\rm tr}({\bf F}_{k}\bar{\boldsymbol{\Phi}})\geq T_{k}(P_{\text{j}}\mathop{\rm tr}({\bf F}_{\text{j}}\bar{\boldsymbol{\Phi}})+$
@@
$\text{s.t.}~p_{k}\mathop{\rm tr}({\bf F}_{k}\bar{\boldsymbol{\Phi}})\geq T_{k}(P_{\text{j}}\mathop{\rm tr}({\bf F}_{\text{j}}\bar{\boldsymbol{\Phi}})+$
@@ IV-C Algorithm Initialization
    $\text{s.t.}~p_{k}^{(0)}\mathop{\rm tr}({\bf F}_{k}\bar{\boldsymbol{\Phi}})\geq T_{k}(P_{\text{j}}\mathop{\rm tr}({\bf F}_{\text{j}}\bar{\boldsymbol{\Phi}})+$
@@ V-A Performance for Various Number of Users with Fixed A-RIS Location
Here we study the performance of the proposed algorithm in terms of the sum total uplink transmit power for $K=2$, $K=3$ and $K=4$ users, for various numbers of jammer antennas and A-RIS elements, assuming a fixed location of the A-RIS. We also assess the benefit and characterize the optimized behavior of the absorption capability of the A-RIS in terms of the average reflection coefficients of the A-RIS elements.
In Table I we show our simulation parameters. Here we assume a fixed network, whereas in Section V-B we relax this by studying the performance for various A-RIS locations. We focus on a scenario with static rather than random channel gains to ensure that near-far relationship between the users remains the same throughout the simulation trials, which (together with the SINR constraints in the optimization) enables us to fix the decoding order rather than adding this as an additional complication to the simulation. The benefit of the A-RIS in mitigating the jamming for the NOMA system is ultimately independent of this assumption, so the purpose is mainly to simplify our experiments. Note that for numerical stability in the optimization, we normalize the noise at the BS as $\sigma^{2}=1$ and scale the path losses correspondingly. The chosen path loss values are representative for a network with inter-node distances of a few tens of meters, assuming path loss exponent $\alpha=3$.
@@
We choose to focus on a jammer with a rather large number of antennas $M$ and also an A-RIS with a large number of elements $N$. This is an interesting case in practice, since in mm-wave systems a rather large number of antennas/elements would be needed, and we want to show the capability of Algorithm 1 to find a solution for such large optimization problems. This is in contrast to our initial work in [25] using the Matlab fmincon tool, in which we could not find an optimal solution for a jammer larger than $M=16$ in combination with A-RIS larger than $N=32$. Note that we assume a maximum A-RIS reflection coefficient $\beta_{\text{max}}=1$, despite our discussions in Section III. The reason is that in our considered evaluation scenarios, a value of $\beta_{\text{max}}<1$ can simply be included in the assumed path losses to the A-RIS.
@@
In Fig. 4 we show the required total user transmit power to meet the user QoS requirements in the presence of a jammer with $M=4,8,16$ antennas, as a function of the number of A-RIS elements $N=32,38,44,64,88,100,128$ for $K=2$ users. As seen, a large A-RIS is very beneficial to mitigate the jammer’s power. The sum transmit power drops from 0.28 W (24.5 dBm) to 0.07 W (18.5 dBm) when increasing the number of A-RIS elements from =4=4 to $N=128$ in the case of a jammer with $M=4$ antenna elements. Doubling the number of jammer antennas to $M=8$ does not significantly change the required total transmit power. However, with $M=16$ there is a significant increase in the required sum user power for A-RIS with $32$ elements, increasing to 0.54 W (27.3 dBm), whereas for large A-RIS the increase in sum user power is still small.
@@
In Fig. 6, we further investigate the absorption behavior of the A-RIS as a function of the number of jammer antennas $M$ and A-RIS elements $N$. Here we investigate larger $M$ and $N$ compared to Fig. 5, but the behavior of the A-RIS is the same. This also shows that Algorithm 1 converges well also for large $M$ and $N$. In general, for the investigated deployment and propagation scenario here, it seems that with $N/M\geq\approx 5$, absorption is not needed, whereas for $N/M<\approx 5$, absorption is very useful. This can be explained by the required number of DoFs for the A-RIS to cancel the jammer signal.
In Fig. 7 and Fig. 8, we investigate the performance and behavior of the A-RIS in the case of $K=3$ users. Algorithm 1 also converges very well in this case, and the overall behavior is the same as in the $K=2$ users case. One difference comparing Fig. 7 with Fig. 4 is that the total required transmit power by the users is larger for $K=3$ compared to $K=2$, due to the residual NOMA multi-user interference for the users $k<K$ that is not canceled at the BS. Comparing Fig. 6 and Fig. 8 shows that the A-RIS absorption decreases for the larger number of users with increased multiuser interference (MUI). This is interesting since absorption increases when the jammer interference increases. The difference can be explained by the fact that the A-RIS is attempting to entirely cancel the jammer, while combating the MUI still requires the individual user signals to be received with a strong SNR.
Figure 7: Required total transmit power to meet the user QoS requirements in the presence of a jammer with $M=4,8,16$ antennas, as a function of the number of A-RIS elements in case of $K=3$ users.
@@
TABLE II: Comparison of required total power in the presence of a jammer with $M=4,8,16$ antennas for $K=2,3,4$ users, with and without A-RIS ($N=32$)
Number of $M$ Number of UE Without A-RIS With A-RIS
$M=4$ 2 Users $\times 10^{3} 0.270.27$
3 Users $0.51\times 10^{4}$ $1.19$
4 Users $1.65\times 10^{4}$ $10.68$
$M=8$ 2 Users $0.59\times 10^{3}$ $0.30$
3 Users $0.89\times 10^{4}$ $1.33$
4 Users $35\times 10^{4} 11.$9011.90
$M=16$ 2 Users $1.07\times 10^{3}$ $0.54$
3 Users $1.90\times 10^{4}$ $3.80$
4 Users $2.72\times 10^{4}$ $36.53$
In Table II we present a comparison of the total required transmit power in the presence of a jammer with different numbers of antennas $M=4,8,16$ for $K=2$, $K=3$ and $K=4$ users, with and without the assistance of an A-RIS. As shown, there is a very large gain with the A-RIS, up to more than a $28$ dB gain in the required total transmit power. The results also clearly show that as the number of jammer antennas $M$ increases, the total power consumption increases more without an A-RIS than with an A-RIS. When increasing the number of users $K$ from $K=3$ to $K=4$, there is a larger increase in the required total transmit power than going from $K=2$ to $K=3$ users, indicating that using the A-RIS degrees of freedom to simultaneously mitigate the jammer and control the multi-user interference is getting more challenging. However, the gain compared to the case with no A-RIS is still substantial.